# Determination of impact parameter for CEE with digi-input neural networks

### v3

@@ 3.2 Regression models
Two DNN models have been implemented for the task: CNN and Graph Attention network (GAT)[39, 40]. The structures of the two models are demonstrated in the Appendix Figure 12 and 13. For both models, a dense layer is implemented at the entrance of the input data to encode the three features for each readout stip. The encoder outputs 762×64 hidden features for further extractions with two sequential CNN or GAT layers. Finally, all the output features are flattened and decoded to the impact parameter prediction by a dense layer.
CNN extracts the correlation between geometrically local features and trainable kernels through convolution, and it is one of the most popular algorithms in image recognition. For this work, the phase space distribution can be treated as an image and is capable of inputs with multi-features, which is similar to the RGB channels of the image. Besides the phase space input, CNN is also capable of processing digitized signal input. In this mode features from the 762 strips are arranged into a 96×7×3 shape to reflect the geometry layout. The convolutional kernel in CNN is a key module for feature extraction, and its size should be modified according to the geometry coverage that the input data includes. The kernel size of the first convolutional layer is set to 5×5 for phase space input, while for digi input it is 3×32 because of the strip dimension. The other hyper-parameters of the models can be found in the Appendix Table 2.
Graph Neural Networks (GNN) have recently risen as a popular machine learning model with the capability of graphic structure handling. For nuclear collisions, data of the produced tracks, hits, or even signals can be regarded as graph data, in which each individual, or node in the concept of a graph, is potentially connected with others under the collision dynamics. GAT is one of the GNNs that provides robustness in dynamic graphs. It trains a universal and trainable agent for calculating the attention weights between any pair of nodes. Then the node features can be transferred as a combination of itself and its neighbouring nodes according to the attention weights. Equation 3.2 and 3.3 describes the principle of the GAT layer [^2] In this work, GATv2 structure[40] is implemented for better reported performance over GAT.. Firstly, the attention weights $\alpha_{ij}$ are calculated, where $\vec{h}_{i}$ stands for the input features of node $i$, $\sigma$ the activation function, $\boldmath W$ and $\vec{a}$ the trainable weight matrix and vector. Then, the output features $\vec{h}_{i}^{\prime}$ is obtained as the attention weighted combination of the encoded input features from the self and neighboring nodes $\mathcal{N}_{i}$.
@@ 3.3 Test of the models
The tests in this paper, along with the key settings, are listed in Table 1. They are designed for different purposes of performance evaluation and validation. Firstly, to evaluate the model performances, CNN, GAT, and fully connected (FC) neural networks are implemented to regress the hit data of eTOF. Secondly, to compare the data inputs, CNN models are trained with phase space data and digitized signal data at the forward angle respectively. Finally, for the discussion of further expansion of acceptance with large angle detectors included, we conduct a comparative study with the CNN model to regress the phase space data of the forward angle and total acceptance of CEE.
In this work, the graph in GAT network is costructed by connecting every two readout strips, which leaves the network to learn the edge attributes. GAT has a masked attention feature which allows eliminating attention weights of some predefined edges. Based on the characteristics, we tested the performances of GAT networks when, for each node, the ones outside the kernel of the CNN (3$\times$32) and the ones within the kernel are masked respectively. The tests, labeled as GAT-HIT-FW-f and GAT-HIT-FW-c, can provide evidences on which nodes, farther or closer, are more important to a specific node.
For all the models, mean square error is used in the training as the loss function. An early-stop strategy is implemented for keeping the best weights of the DNNs when the performance in the valid data is not improved for 10 epochs. The best weights were reached after 10-40 epochs of training, depending on the models. Overfitting is not visible in this work since the data amounts are large enough to represent the overall feature distribution. Moreover, the tests have been evaluated on their stability with different train-test data splits.
@@
[…]
@@ 4 Performances and discussion
The performance of the tests is evaluated by the following indicators: Mean Absolute Error (MAE), Mean Square Error (MSE) and goodness of fit $R^{2}$. MAE and MSE are calculated with Equation 4.1 and 4.2 respectively:
@@
where $n$ is the size of the data for validation or test, and $\overline{b}$ stands for the average value of the impact parameters.
@@
Table 1 lists the performances of the tests under the $b{\rm d}b$ test set. It is reasonable to find that the CNN-PS-TOT model has the best performance for the input features from the total phase space. It can be treated as an estimation of the determination capability in the condition of the CEE spectrometer and energy. The performance of CNN-PS-FW shows the capability with phase space identified by the forward detectors, and it also serves as a reference model for comparison with others that take the forward TOF data as input. It is inspiring to find that the models with eTOF hit input show superior performances than the reference model, with decreases of the MAE by 6-18%. Such an improvement in the HIT-FW tests indicates that the digitized signal data, despite the ambiguity, preserves implicit patterns that can be extracted by DNN models in diverse forms. For example, the position information can potentially increase the equivalent granularity of the HIT-FW models, and the time may reveal the secondary particle types. Yet, these features are no longer reserved in the phase space inputs.
@@
Among the HIT-FW tests, the GAT model has the highest precision in centrality prediction. The dependency of the prediction resolution (MAE) and impact parameters for the HIT-FW tests is investigated, as shown in Figure 7. Each data point is defined on the uniform test set with an impact parameter interval of [-0.25, 0.25) fm. For example, the first point corresponds to the data in [0, 0.5) fm. The three models show some similar patterns in their curves but for most centrality including the central region, the GAT model outperforms other models significantly given the uncertainty of 0.4% (50k events for each data point).
Figure 7: (a) MAE deviations of digi-input models with respect to impact parameters. (b) Relative MAE deviations in the function of impact parameters.
@@
The aggregation mechanism of neighbors in the GAT structure contributes to the overall superiority because it helps eliminate the randomness of the single hits. Such a principle is also embodied in the CNN model which is based on spatial convolution. However, the size of the convolution kernel confines the range of the aggregation locally. It can be seen that masking either the local (GAT-HIT-FW-f) or the distant nodes (GAT-HIT-FW-c) from transfering their features will decrease the prediction power of the GAT model.
Figure 8: (a) Biases of digi-input models with respect to impact parameters. (b) Relative biases of digi-input models with respect to impact parameters.
@@
Figure 8 shows the relative prediction biases of the HIT-FW tests with centrality. The three models show satisfying bias suppression. For very central collisions (b<1.5 fm), the relative bias rises significantly, likely due to the high prediction uncertainty in this centrality region. The least accuracy of the GAT model will not have significant influence on its application, given the best resolution which includes the bias.
@@
The prediction of the GAT model on the $b{\rm d}b$ test dataset is examined visually in Figure 9. A good correlation is shown which agrees with the results in Table 1. It seems that the linear relationship is interfered at around 6 fm, where the prediction error is higher than other ranges in Figure 7. Similar patterns are visible for other models as well. Figure 10 shows the dependency of impact parameters to the correlated collision features. $b_{\rm CNN-TOT}$, $b_{\rm CNN-FW}$, $b_{\rm GAT}$ represent the predicted impact parameters from model CNN-PS-TOT, CNN-PS-FW and GAT-HIT-FW respectively, $M_{\rm tot}$ the multiplicity of total acceptance, $p_{\rm T,tot}$ the summation of reduced transverse momentum for charged particles in the overall accepatance, $M_{\rm fw}$, $p_{\rm T,fw}$ the ones in the forward acceptance. It is shown that the correlation of the features at forward acceptance is weak for impact parameters around 6 fm. It may explain the low precision of such a region for forward predictors. When the impact parameter decreases, the secondary particle distribution moves toward the mid-rapidity region which goes beyond the coverage of eTOF. Consequently, the forward quantities lose the monotonicity and decline, which supports why the model fed with overall phase spaces shows the best performance. In general, the models trained from either the phase spaces or the digitized signals succeed in reproducing the dependencies between the impact parameter and the collision quantities as simulation shows.
@@
Figure 10: Dependency of the true and predicted impact parameters to the correlated collision features.
A significant characteristic of the GAT model is the interpretability of graphical relations. We investigate the importance of the attention weights among the 672 readout strips with the help of NetworkX[41], a tool for visualizing and analysing the graph. Figure 11 shows the graph visualization of the GAT model in predicting the test dataset. The grey cloud is composed of an intense crowd of black edges whose widths stand for the average value of the attention weights. The nodes in the graph are arranged following the Fruchterman-Reingold force-directed algorithm[42], and as a result, nodes with stronger connections to others will be dragged to the central area. The nodes in the graph are colored according to the total signal counts on the strip. It is observed that the light-colored nodes, which represent the strips from the inner area of eTOF, have a higher importance in the contribution of the features, which indicates that the features from these strips can be effectively transmitted to the other strips, even for those at the outer region.
@@ 5 Conclusion
This work is one of the efforts on the determination of impact parameters for CEE. DNN-based models are regarded as appropriate approaches as high dimensional features are needed in low collision energies. We propose that the original digitized signals from the eTOF MRPCs are informative as input, and the DNN models are powerful enough to extract the predictive features. Several structures of DNN models are implemented and tested, showing that the prediction precision with digitized signal inputs exceeds the phase space inputs of the same acceptance by up to 18%. Among the models in our tests, the GAT model reaches the best prediction precision of 0.479 fm in mean absolute error, and the prediction power outperforms other models significantly for central collisions. It is also examined that larger acceptance will result in better performance, which encourages further investigations of including the channel responses from other subsystems.
@@ Appendix A Structures and hyper-parameters of the models
Figure 13: Structure of the CNN network used in this work. The upper flow processes digitized signal inputs while the lower flow processes phase space inputs.
@@
Learning rate 5x10-4 Learning rate 5x10-4
@@
Learning rate 5x10-4 Learning rate 5x10-4

### v4

@@ 3.2 Regression models
Two DNN-based models have been implemented: CNN and Graph Attention network (GAT)[39, 40]. The structures of the two models are demonstrated in the Appendix Figure 12 and 13. For both models, a dense layer is implemented at the entrance of the input data to encode the three features for each readout strip (762$\times$3). The encoder outputs 762$\times$64 hidden features for further extractions with two sequential CNN or GAT layers. A final dense layer flattens all the output features and decodes them to the impact parameter prediction.
CNN extracts the correlation between geometrically local features and trainable kernels through convolution, and it is one of the most popular algorithms in image recognition. For this work, the phase space distribution can be treated as an image and is capable of inputs with multi-features, similar to the RGB channels of images. Besides the phase space input, CNN can also process digitized signal input. In this mode, features from the 762 strips are arranged into a 96$\times$7$\times$3 shape to reflect the geometry layout. The convolutional kernel in CNN is a crucial module for feature extraction, and its size should be modified according to the geometry coverage of the input data. The kernel size of the first convolutional layer is set to 5$\times$5 for phase space input, while for digi input it is 32$\times$3 because of the strip dimension. The other hyper-parameters of the models can be found in the Appendix Table 2.
Graph Neural Networks (GNN) have recently become a popular machine learning model that can handle graphic structures. For nuclear collisions, data of the produced tracks, hits, or even signals can be regarded as graphs, in which each individual, or node in the concept of a graph, is potentially connected with others under the collision dynamics. Implementing GNNs through the GAT architecture enables robustness in dynamic graphs. It trains a universal and trainable agent for calculating the attention weights between any pair of nodes. Then, the node features can be transferred as a combination of themselves and those of its neighboring nodes, according to the attention weights. Equation 3.2 and 3.3 describes the principle of the GAT layer [^3] In this work, GATv2 structure[40] is implemented for better reported performance over GAT.. Firstly, the attention weights $\alpha_{ij}$ are calculated, where $\vec{h}_{i}$ stands for the input features of node $i$, $\sigma$ the activation function, $\boldmath W$ and $\vec{a}$ the trainable weight matrix and vector. Then, the output features $\vec{h}_{i}^{\prime}$ is obtained as the attention-weighted combination of the encoded input features from the self and neighboring nodes $\mathcal{N}_{i}$.
@@ 3.3 Test of the models
The tests described in this paper, along with their key settings are listed in Table 1. They are designed for different purposes of performance evaluation and validation. Firstly, to evaluate the model performances, CNN, GAT, and fully connected (FC) neural networks are implemented to regress the hit data of eTOF. Secondly, to compare the input types, CNN models are trained with phase space data and digitized signal data at the forward angle. Finally, for the discussion of further expansion of acceptance with large angle detectors included, we conduct a comparative study with the CNN model to regress the phase space data of the forward angle and total acceptance of CEE.
In this work, the graph in the GAT network is constructed by connecting every two readout strips, which leaves the network to learn the edge attributes. GAT has a masked attention feature that eliminates attention weights of some predefined edges. Based on the characteristics, we tested the performances of GAT networks when, for each node, the ones outside the kernel of the CNN (32$\times$3) and the ones within the kernel are masked respectively. The tests, labeled as GAT-HIT-FW-f and GAT-HIT-FW-c, can provide evidence on which nodes, farther or closer, are more important to a specific node.
For all the models, mean square error is used as the loss function in training. An early-stop strategy is implemented to keep the best weights of the DNNs when the performance in the valid data is not improved for 10 epochs. The best weights were reached after 10-40 training epochs, depending on the model. Overfitting is not visible in this work since the data amounts are large enough to represent the overall feature distribution. Moreover, the tests have been evaluated on their stability with different train-test data splits.
@@
Test label Acceptance Type of data input Network MAE [fm] MSE [fm2] $R^{2}$
@@ 4 Performances and discussion
The following indicators evaluate the performances of the models: Mean Absolute Error (MAE), Mean Square Error (MSE), and goodness of fit $R^{2}$. MAE and MSE are calculated with Equation 4.1 and 4.2 respectively:
@@
where $n$ is the size of the data for validation or test and $\overline{b}$ stands for the average value of the impact parameters.
@@
Table 1 lists the performances of the models under the $b{\rm d}b$ test set. It is reasonable to find that the CNN-PS-TOT model has the best performance for the input features from the total phase space. It can be treated as an estimation of the determination capability in the condition of the CEE spectrometer and energy. The performance of CNN-PS-FW shows the capability with phase space identified by the forward detectors, and it also serves as a reference model for comparison with others that take the forward TOF data as input. It is inspiring to find that the models with eTOF hit input perform better than the reference model, with decreases of the MAE by 6-18%. Such an improvement in the HIT-FW tests indicates that the digitized signal data preserve implicit patterns that can be extracted by DNN models in diverse forms. For example, the position information can potentially increase the equivalent granularity of the HIT-FW models, and the time may reveal the secondary particle types. Yet, these features are no longer preserved in the phase space inputs.
With phase space data and CNN-based models, CNN-PS-TOT exhibits superior precision compared to CNN-PS-FW. Notably, this precision improvement, attributed to acceptance, is more significant than data format and DNN model. This phenomena indicates that the prediction uncertainty of the model GAT-HIT-FW is primarily aleatoric rather than epistemic. Further quantitative assessment of the uncertainty sources is not covered in this work which only gives error definitions based on ground truth.
@@
Among the HIT-FW tests, the GAT model has the highest precision in centrality prediction. The dependency of the prediction resolution (MAE) on the impact parameter for the HIT-FW tests is investigated, as shown in Figure 7. Each data point is defined on the uniform test set with an impact parameter interval of [-0.25, 0.25) fm. For example, the first point, at 0.25 fm, corresponds to the data in [0, 0.5) fm. The three models show similar patterns in their curves. Still, for most centrality, including the central region, the GAT model outperforms other models significantly, given the statistical uncertainty of 0.4% (50k events for each data point).
Figure 7: Prediction MAE as a function of impact parameter.
@@
The aggregation mechanism of neighbors in the GAT structure contributes to the overall superiority because it helps eliminate the randomness of the single hits. Such a principle is also embodied in the CNN model, which is based on spatial convolution. However, the convolution kernel size confines the aggregation range. It can be seen that masking either the local (GAT-HIT-FW-f) or the distant nodes (GAT-HIT-FW-c) from transferring their features will decrease the prediction power of the GAT model.
Figure 8: Prediction Bias as a function of the impact parameter.
@@
Figure 8 shows the relative prediction biases of the HIT-FW tests as a function of centrality. The three models show satisfying bias suppression. For very central collisions (b<1.5 fm), the relative bias rises significantly. The lowest accuracy of the GAT model will not significantly influence its application, given the best prediction MAE which has taken the bias into account.
@@
The prediction of the GAT model on the $b{\rm d}b$ test dataset is examined visually in Figure 9. A good correlation is shown, which agrees with the results in Table 1. The linear relationship seems to be lost at around 6 fm, where the prediction error is higher than in other ranges in Figure 7. Similar patterns are visible for other models as well. Figure 10 shows the dependency of impact parameters on the correlated collision features. $b_{\rm CNN-TOT}$, $b_{\rm CNN-FW}$, $b_{\rm GAT}$ represent the predicted impact parameters from model CNN-PS-TOT, CNN-PS-FW and GAT-HIT-FW respectively, $M_{\rm tot}$ the multiplicity of total acceptance, $p_{\rm T,tot}$ the summation of reduced transverse momentum for charged particles in the overall acceptance, $M_{\rm fw}$, $p_{\rm T,fw}$ the ones in the forward acceptance. It is shown that the correlation of the features at forward acceptance is weak for impact parameters around 6 fm. This may explain the low precision of such a region for forward predictors. When the impact parameter decreases, the secondary particle distribution moves toward the mid-rapidity region, which goes beyond the coverage of eTOF. Consequently, the forward quantities lose the monotonicity and decline, which supports why the model fed with overall phase spaces performs best. In general, the models trained from either the phase spaces or the digitized signals successfully reproduce the correlations between the impact parameter and the collision quantities.
@@
Figure 10: Dependency of the true and predicted impact parameters on the correlated collision features.
A significant characteristic of the GAT model is the interpretability of graphical relations. We investigate the importance of the attention weights among the 672 readout strips with the help of NetworkX[41], a tool for visualizing and analyzing the graph. Figure 11 shows the graph visualization of the GAT model in predicting the test dataset. The grey cloud includes an intense crowd of black edges whose widths represent the average value of the attention weights. The nodes in the graph are arranged following the Fruchterman-Reingold force-directed algorithm[42], and as a result, nodes with stronger connections to others will be dragged to the central area. The nodes in the graph are colored according to the total signal counts on the strip. It is observed that the light-colored nodes, which represent the strips from the inner area of eTOF, have a higher importance in the contribution of the features, which indicates that the features from these strips can be effectively transmitted to the other strips, even to those at the outer region.
@@ 5 Conclusion
This work is one of the efforts to determine the impact parameter for CEE. DNN-based models are regarded as appropriate approaches as high dimensional features are needed in low collision energies. We propose that the original digitized signals from the eTOF MRPCs are informative as input and that the DNN models are powerful enough to extract the predictive features. Several structures of DNN models are implemented and tested, showing that the prediction precision with digitized signal inputs exceeds the phase space inputs of the same acceptance by up to 18%. Among the models in our tests, the GAT model reaches the best prediction precision of 0.479 fm in mean absolute error, and the prediction power outperforms other models significantly for central collisions. It is also examined that larger acceptances will result in better performance, which encourages further investigations of including data from other subsystems.
Acknowledgments
@@ Appendix A Structures and hyper-parameters of the models
Figure 13: Structure of the CNN network used in this work. The upper flow processes digitized signal inputs, while the lower flow processes phase space inputs.
@@
Learning rate 5$\times$10-4 Learning rate 5$\times$10-4
@@
Learning rate 5$\times$10-4 Learning rate 5$\times$10-4